An investment project requires a net investment of $200 million. The project is expected to generate annual net cash flows of $25 million for the next 15 years with a one-time end of project cash flow of $3 million. The firm's cost of capital is 14 percent and marginal tax rate is 40 percent.

a) Evaluate the project using the NPV method and state whether or not the project should be accepted.

b) Evaluate the project using the IRR method and state whether or not the project should be accepted.

c) Evaluate the project using the PI method and state whether or not the project should be accepted.

NPV = -$46 million < 0 so Reject; IRR = 9.20 < 14% so Reject; PI = 0.77 < 1 so Reject

Show formulas used
Answer #1

a). NPV = -initial investment + PV of annual net cash flows for 15 years + PV of after-tax project end cash flow

= -200 + 25*PVIFA(14%,15) + 3/(1+14%)^15 = -46 million.

NPV is negative so project should not be accepted as it will be loss making.

b). IRR is calculated the IRR function with the given cash flows: CF0 = -200; CF1 to CF14 = 25; CF15 = 25 + 3 = 18

Solving, IRR = 9.20%

The project should not be accepted since IRR is less than the firm's cost of capital.

c). Profitability index = PV of future cash flows/initial investment

= [25*PVIFA(14%,15) + 3/(1+14%)^15]/200 = 0.77

The project should not be accepted as it has a profitability index of less than 1 indicating that it will not recover its initial investment.